# Label Words are Anchors: An Information Flow Perspective for Understanding In-Context Learning

### v1

@@ Abstract
In-context learning (ICL) emerges as a promising capability of large language models (LLMs) by providing them with demonstration examples to perform diverse tasks. However, the underlying mechanism of how LLMs learn from the provided context remains under-explored. In this paper, we investigate the working mechanism of ICL through an information flow lens. Our findings reveal that label words in the demonstration examples function as anchors: (1) semantic information aggregates into label word representations during the shallow computation layers’ processing; (2) the consolidated information in label words serves as a reference for LLMs’ final predictions. Based on these insights, we introduce an anchor re-weighting method to improve ICL performance, a demonstration compression technique to expedite inference, and an analysis framework for diagnosing ICL errors in GPT2-XL. The promising applications of our findings again validate the uncovered ICL working mechanism and pave the way for future studies.
@@ 1 Introduction
In-context Learning (ICL) has emerged as a powerful capability alongside the development of scaled-up large language models (LLMs) (Brown et al., 2020). By instructing LLMs using few-shot demonstration examples, ICL enables them to perform a wide range of tasks, such as text classification (Min et al., 2022a) and mathematical reasoning (Wei et al., 2022). Since ICL does not require updates to millions or trillions of model parameters and relies on human-understandable natural language instructions (Dong et al., 2023), it has become a promising approach for harnessing the full potentiality of LLMs. Despite its significance, the inner working mechanism of ICL remains an open question, garnering considerable interest from research communities (Xie et al., 2022; von Oswald et al., 2022; Dai et al., 2022; Olsson et al., 2022).
@@
Figure 1: Saliency visualization results of shallow and deep layers of a GPT model. Here, the depth of the line from the right word to the left word reflects the importance of the information flow in ICL.
@@
In this paper, we find that the label words serve as anchors that aggregate and distribute information in ICL. Specifically, we first visualize the attention interactive pattern between tokens with a GPT model (Brown et al., 2020) on sentiment analysis. As shown in Figure 1, we have an intuitive observation that as the layer goes deeper, the label words in the demonstration will become more dominant for the prediction. To draw a clearer picture of this phenomenon, we compute two metrics based on saliency scores to portray the information flow in ICL and further propose the following hypothesis: Information Flow with Labels as Anchors $\mathcal{H}_{1}$: In shallow layers, label words gather the information of demonstrations to form semantic representations for deeper layers. $\mathcal{H}_{2}$: In deep layers, the model extracts the information from label words to form the final prediction.
@@
We design two experiments to validate the hypothesis using GPT2-XL (Radford et al., 2019) and GPT-J (Wang and Komatsuzaki, 2021) on various text classification benchmarks. (1) By isolating the label words in certain layers to block the information aggregation path to the label words, we find that such isolation in shallow layers significantly impairs model performance. This indicates that label words indeed collect useful semantics during the forward propagation in shallow layers. (2) We examine the correlation between the attention distributions on the label words of the target position and the final prediction results. The results show that the prediction positively correlates with the attention weights on label words, i.e., the probability of a candidate is higher with more attention weights on the specific label. In summary, these experimental findings suggest that our hypothesis holds well with large language models on real-world datasets.
@@
Drawing on insights from the information flow perspective, we explore three approaches to enhance ICL’s effectiveness, efficiency, and interpretability. (1) We introduce an anchor re-weighting method utilizing a learnable vector to adaptively adjust the significance of various label words in demonstration examples, achieving a 16.7 average accuracy improvement over vanilla ICL baselines. (2) To expedite ICL inference, we compress its input into pre-computed anchor representations, as model predictions primarily rely on label word activations. Experiments demonstrate that the inference can be accelerated 1.8 $\times$ with negligible performance degradation. (3) We present an error analysis example using ICL on GPT2-XL, revealing that the label confusion matrix closely mirrors the distance distribution of anchor key vectors, suggesting errors may arise from indistinguishable anchor representations. These promising applications further validate our hypothesis and shed light on future ICL studies for better transparency of LLMs.
@@ 2 Label Words are Anchors
In this section, we first confirm our intuitive findings with two metrics based on saliency scores in § 2.1. Based on the quantitative results, we further propose the hypothesis to interpret the working mechanism of ICL: $\mathcal{H}_{1}$: In shallow layers, label words aggregate information from demonstration examples to form semantic representations for later computations. $\mathcal{H}_{2}$: In deep layers, the model makes predictions by extracting information from label words. We validate these two hypothesis in § 2.2 and § 2.3, respectively.
@@ 2.1 Hypothesis Motivated by Saliency Scores
In this part, we aim to visualize the attention interactive pattern between tokens for GPT2-XL, and find patterns behind such interaction. We utilize a common interpretation tool, saliency technique Simonyan et al. (2013), to reveal the important token interactions. Following common practice, we use the Taylor expansion Michel et al. (2019) to calculate the saliency score for each element of the attention matrix:
$I_{l}=\sum_{h}\left|A_{h,l}^{\top}\frac{\partial\mathcal{L}(x)}{\partial A_{h,l}}\right|$ (1)
Here, $A_{h,l}$ is the value of the attention matrix of the $h$-th attention head in the $l$-th layer, $x$ is the input, and $\mathcal{L}(x)$ is the loss function of the task, e.g., the cross-entropy objective for a classification problem. We average all attention heads to obtain the saliency matrix $I_{l}$ for the $l$-th layer. $I_{l}(i,j)$ represents the importance of the information flow from the $j$-th word to the $i$-th word for ICL. By observing $I_{l}$, we can get an intuitive impression that as the layer goes deeper, demonstration label words will become more dominant for the prediction, as depicted in Figure 1.
To draw a clearer picture of this phenomenon, we propose the following quantitative metrics base on $I_{l}$. Let the positions of the label words in the input $x$ be $p_{1},...,p_{C}$, the target position be $q$, and the saliency matrix of the $l$-th layer be $I_{l}$. Then, we define three quantitative metrics:
$\boldsymbol{S}_{wp}$, the average importance of the information flow from other words to label words:
@@
$\boldsymbol{S}_{pq}$, the average importance of the information flow from label words to the target position:
@@
$C_{pq}$ $=\{(p_{k},q):k\in[1,C]\}.$
$\boldsymbol{S}_{ww}$, the average importance of the information flow between any words (excluding $S_{wp}$ and $S_{pq}$):
@@
$C_{ww}=$ $\{(i,j):j<i\}$
$-C_{wp}-C_{pq}.$
@@
We choose GPT2-XL from the GPT series Radford et al. (2019) as the main model to be examined, due to its moderate model size (of 1.5B parameters) that is suitable for our hardware resource and its decent ICL performance (Dai et al., 2022). For datasets, we use a sentiment analysis task, Stanford Sentiment Treebank Binary (SST-2) Socher et al. (2013), a question type classification task, Text REtrieval Conference Question Classification (TREC) Li and Roth (2002); Hovy et al. (2001), a topic classification task, AG’s news topic classification dataset (AGNews) Zhang et al. (2015), an emotion classification task, EmoContext (EmoC) Chatterjee et al. (2019). The templates used for constructing demonstrations of these datasets can be found in Appendix A. We extracted $1000$ examples from the test set for evaluation. For ease of analysis, we sample one example for each class from the training set to form the demonstration, and the demonstration order follows a random order. All results are averaged over five random seeds.
@@
Figure 2: Relationship between the relative sizes of $S_{wp}$, $S_{pq}$, and $S_{ww}$ and the number of layers. (We normalized $(S_{wp},S_{pq},S_{ww})$) Results of TREC and EmoC can be seen in Apendix D. Initially, $S_{wp}$ occupies a significant proportion, but it gradually decays over the layers, while $S_{pq}$ becomes the dominant one.
The results are shown in Figure 2. It can be seen that (1) in the shallow layers, $S_{pq}$, the importance of the information flow from label words to targeted positions is low, while $S_{wp}$, the information flow from other words to label words is high; (2) in the deep layers, $S_{pq}$, the importance of information flow from label words to the targeted position become the dominant one. Meanwhile, $S_{pq}$ and $S_{wp}$ are usually more than $S_{ww}$, indicating that the interactions involving label words are more important than others.
@@
Based on this, we propose the hypothesis that label words function as anchors in the ICL information flow. In shallow layers, label words gather information from demonstration examples to form semantic representations for deeper layers, while in deep layers, the model extracts the information from label words to form the final prediction.
@@ 2.2 Shallow Layers: Information Aggregation
In this part, we validate the first part of our hypothesis. We hypothesize that the aggregation of information in in-context learning relies on the flow of information from demonstration tokens to label tokens, which is facilitated by the attention mechanism in transformers. By manipulating attention to block this information flow and examining the changes in model behavior, we can validate whether the information aggregation process exists and how much it contributes to the final predictions.
@@
We maintain the same sample size of 1000 inputs from the test set as Section 2.1. We use the same demonstration for a single random seed. To further validate our findings on larger models, we incorporate GPT-J (6B) (Wang and Komatsuzaki, 2021) in our experiments, which surpasses GPT2-XL in terms of model size and capacity.
@@
To achieve the intervention on the information flow on the label words, we isolate label words in certain layers, by disabling the attention from the label words attending to the demonstrations. Specifically, we set $A_{l,h}(p,i)(i<p)$ in the attention matrix $A_{l,h}$ of each attention head in the $l$-th layer to 0, where $p$ is the position of the label word, and $i$ is the position of the word before the label word. Therefore, label words in the $l$-th layer cannot receive the information from previous demonstration examples via the attention mechanism.
@@
To measure the effects of blocking the information flow from demonstration tokens to label tokens, we use the following metrics to examine the changes in the model predictions: (1) Label Loyalty: which measures the consistency between the outputs labels with and without isolation. Higher label loyalty indicates that the model predictions are less affected by the isolation. (2) Word Loyalty: which probes more fine-grained changes on the model predictions. Specifically, we adopt the Jaccard similarity between the top-$5$ potential words drawn from the original vocabulary distributions and that after isolation. This metric could capture subtle changes in the output distributions, where higher word loyalty indicates better consistency between the language model outputs. We refer readers to Appendix B for a detailed discussion.
Isolation Layer Label Loyalty Word Loyalty
GPT2-XL No isolation 100.00 100.00
   First 5 layers 44.03 6.30
   Last 5 layers 99.61 99.52
GPT-J No isolation 100.00 100.00
   First 5 layers 62.13 53.77
   Last 5 layers 99.01 97.76
Table 1: Effects of isolating label words, results are averaged across the SST-2, TREC, AGNews, and EmoC datasets (percentage). Isolating the first 5 layers significantly reduces the loyalty.
@@
We isolate the label words in the first 5 layers to examine the existence and effect of information aggregation. We also isolate the label words in the last 5 layers for comparison. As shown in Table 1, isolating label words in the first $5$ layers can cause serious interference to the model’s task completion, while the impact of the last $5$ layers is small. This verifies the existence of information aggregation in shallow layers, and demonstrates its significance in ICL. We also explore the isolation effects by varying the number of labels disabled in Appendix C and observe a similar trend.
@@ 2.3 Deep Layers: Information Extraction
We further validate the second part of our hypothesis, i.e., the model extracts the information from label words to form the final prediction. We examine the correlation between the attention distributions on the label words of the target position and the model’s final prediction.
We use the same dataset and model setup as in § 2.2.
@@ 2.3.1 Experiments
Suppose the positions of the label words in the input $x$ are $p_{1},...,p_{C}$}, the targeted position is $q$, and the sum of the attention matrices of each attention head at the $l$-th layer is $A_{l}$. We postulate that there’s a strong correlation between the attention distributions on the label words of the target position $(A_{l}(q,p_{1}),...,A_{l}(q,p_{C}))$ and the model’s final prediction. We use the AUC-ROC score to quantify this correlation, which we denote as $\text{AUCROC}_{l}$.
We use AUC-ROC for two reasons: (1) Considering the attention mechanism, the attention values are used to weigh the key vectors. The size of the attention cannot fully reflect the importance of the corresponding word; it must be combined with factors such as the norm of the key vector Kobayashi et al. (2020). The AUC-ROC metric can implicitly consider these factors to better discover the correlation. (2) The proportion of different labels output by the model may be unbalanced. Using AUC-ROC can to some extent alleviate this problem and prevent the influence of class imbalance from disturbing our analysis.
Considering the residual mechanism that is common in Transformer models, the hidden state of each layer can be seen as the accumulation of the hidden states calculated separately by the previous layers. To measure the contribution of all layers up to the $l$th layer, we define the accumulated AUC-ROC score $R_{l}$:
@@
Here, we quantify the positive contribution by calculating the difference between AUC-ROC and a baseline threshold of $0.5$. The value of $R_{l}$ represents the ratio of the contribution of the attention distributions on the label words of the target position in all layers up to the $l$-th layer.

### v2

@@ Abstract
In-context learning (ICL) emerges as a promising capability of large language models (LLMs) by providing them with demonstration examples to perform diverse tasks. However, the underlying mechanism of how LLMs learn from the provided context remains under-explored. In this paper, we investigate the working mechanism of ICL through an information flow lens. Our findings reveal that label words in the demonstration examples function as anchors: (1) semantic information aggregates into label word representations during the shallow computation layers’ processing; (2) the consolidated information in label words serves as a reference for LLMs’ final predictions. Based on these insights, we introduce an anchor re-weighting method to improve ICL performance, a demonstration compression technique to expedite inference, and an analysis framework for diagnosing ICL errors in GPT2-XL. The promising applications of our findings again validate the uncovered ICL working mechanism and pave the way for future studies. [^1] https://github.com/lancopku/label-words-are-anchors
@@ 1 Introduction
In-context Learning (ICL) has emerged as a powerful capability alongside the development of scaled-up large language models (LLMs) (Brown et al., 2020). By instructing LLMs using few-shot demonstration examples, ICL enables them to perform a wide range of tasks, such as text classification (Min et al., 2022a) and mathematical reasoning (Wei et al., 2022). Since ICL does not require updates to millions or trillions of model parameters and relies on human-understandable natural language instructions (Dong et al., 2023), it has become a promising approach for harnessing the full potentiality of LLMs. Despite its significance, the inner working mechanism of ICL remains an open question, garnering considerable interest from research communities (Xie et al., 2022; Dai et al., 2022; Akyürek et al., 2022; Li et al., 2023b).
@@
Figure 1: Visualization of the information flow in a GPT model performing ICL. The line depth reflects the significance of the information flow from the right word to the left. The flows involving label words are highlighted. Label words gather information from demonstrations in shallow layers, which is then extracted in deep layers for final prediction.
Figure 2: Illustration of our hypothesis. In shallow layers, label words gather information from demonstrations to form semantic representations for deeper processing, while deep layers extract and utilize this information from label words to formulate the final prediction.
@@
In this paper, we find that the label words serve as anchors that aggregate and distribute information in ICL. We first visualize the attention interactive pattern between tokens with a GPT model (Brown et al., 2020) on sentiment analysis (Figure 1). Initial observations suggest that label words aggregate information in shallow layers and distribute it in deep layers. [^2] In this paper, “shallow” or “first” layers refer to those closer to the input, while “deep” or “last” layers are closer to the output. Here, “deep layers” include those around the midpoint, e.g., layers 25-48 in a 48-layer GPT2-XL. To draw a clearer picture of this phenomenon, we design two metrics based on saliency scores to portray the information flow in ICL and further propose the following hypothesis: Information Flow with Labels as Anchors $\mathcal{H}_{1}$: In shallow layers, label words gather the information of demonstrations to form semantic representations for deeper layers. $\mathcal{H}_{2}$: In deep layers, the model extracts the information from label words to form the final prediction.
@@
Two experiments are designed to validate the hypothesis using GPT2-XL (Radford et al., 2019) and GPT-J (Wang and Komatsuzaki, 2021) across several text classification benchmarks. (1) By blocking the information aggregation path to label words in certain layers, we find that such isolation in shallow layers significantly impairs model performance. This indicates that label words collect useful information during forward propagation in shallow layers. (2) We investigate the relationship between the attention distributions on the label words of the target position and the model’s final prediction. Our results illustrate a strong positive correlation, where a candidate label’s probability increases with more attention weight on its corresponding label token. In summary, these experimental findings suggest that our hypothesis holds well with large language models on real-world datasets.
@@
Drawing on insights from the information flow perspective, we explore three approaches to enhance ICL’s effectiveness, efficiency, and interpretability. (1) An anchor re-weighting method is introduced, which employs a learnable vector to adjust the significance of different label words in demonstrations, leading to a 16.7% average accuracy boost compared to standard ICL baselines. (2) For quicker ICL inference, inputs are compressed into pre-calculated anchor representations since model predictions primarily rely on label word activations. Testing shows a 1.8 $\times$ speedup in inference with only a minimal performance trade-off. (3) An error analysis of ICL on GPT2-XL demonstrates that the label confusion matrix aligns closely with the distance distribution of anchor key vectors, implying that errors might result from similar anchor representations. These promising applications further validate our hypothesis and shed light on future ICL studies for better transparency of LLMs.
@@ 2 Label Words are Anchors
This section confirms the intuitive findings using two saliency score-based metrics as discussed in § 2.1. The quantitative results lead to a proposed hypothesis for the ICL working mechanism: $\mathcal{H}_{1}$: In shallow layers, label words aggregate information from demonstration examples to form semantic representations for later computations. $\mathcal{H}_{2}$: In deep layers, the model makes predictions by extracting information from label words. The validation for these hypotheses is presented in § 2.2 and § 2.3, respectively.
@@ 2.1 Hypothesis Motivated by Saliency Scores
This section aims to discover the inherent patterns in the attention interaction between tokens for a GPT model. The saliency technique Simonyan et al. (2013), a common interpretation tool, is employed for highlighting critical token interactions. Following common practice, we use the Taylor expansion Michel et al. (2019) to calculate the saliency score for each element of the attention matrix:
$I_{l}=\sum_{h}\left|A_{h,l}^{\top}\frac{\partial\mathcal{L}(x)}{\partial A_{h,l}}\right|.$ (1)
Here, $A_{h,l}$ is the value of the attention matrix of the $h$-th attention head in the $l$-th layer, $x$ is the input, and $\mathcal{L}(x)$ is the loss function of the task, e.g., the cross-entropy objective for a classification problem. We average all attention heads to obtain the saliency matrix $I_{l}$ for the $l$-th layer. $I_{l}(i,j)$ represents the significance of the information flow from the $j$-th word to the $i$-th word for ICL. By observing $I_{l}$, we can get an intuitive impression that as the layer goes deeper, demonstration label words will become more dominant for the prediction, as depicted in Figure 1.
To draw a clearer picture of this phenomenon, we propose three quantitative metrics based on $I_{l}$. Our focus lies in three components: (i) the label words, such as “Negative” and “Positive” in Figure 2, denoted as $p_{1},...,p_{C}$, where $C$ represents the total number of label words; [^3] In this study, the term ’label words’ is approximately equal to ’label tokens’. The only deviation is the ’Abbreviation’ in the TREC dataset, where we use the first subword in experiments, following Zhao et al. (2021). (ii) the target position, where the model generates prediction labels (i.e., the final token in the input), which we denote as $q$; and (iii) the text part, i.e., the tokens before label words in the demonstration.
The definitions of the three quantitative metrics follow below.
$\boldsymbol{S}_{wp}$, the mean significance of information flow from the text part to label words:
@@
$\boldsymbol{S}_{pq}$, the mean significance of information flow from label words to the target position:
@@
$C_{pq}$ $=\{(q,p_{k}):k\in[1,C]\}.$
$\boldsymbol{S}_{ww}$, the mean significance of the information flow amongst all words, excluding influences represented by $\boldsymbol{S}_{wp}$ and $\boldsymbol{S}_{pq}$ :
@@
$C_{ww}=$ $\{(i,j):j<i\}-C_{wp}-C_{pq}.$
$S_{wp}$, $S_{pq}$, and $S_{ww}$ help assess different information flows in the model. $S_{wp}$ indicates the intensity of information aggregation onto label words. A high $S_{pq}$ demonstrates a strong information extraction from label words for final decision-making. $S_{ww}$ assesses average information flow among words, serving as a benchmark to gauge the intensity of the patterns identified by $S_{wp}$ and $S_{pq}$.
@@
We choose GPT2-XL from the GPT series Radford et al. (2019) as our primary model for investigation, due to its moderate model size (of 1.5B parameters) that is suitable for our hardware resource and its decent ICL performance (Dai et al., 2022). For datasets, we use Stanford Sentiment Treebank Binary (SST-2) Socher et al. (2013) for sentiment analysis, Text REtrieval Conference Question Classification (TREC) Li and Roth (2002); Hovy et al. (2001) for question type classification, AG’s news topic classification dataset (AGNews) Zhang et al. (2015) for topic classification, and EmoContext (EmoC) Chatterjee et al. (2019) for emotion classification. Templates for constructing demonstrations are provided in Appendix A. $1000$ examples are sampled from the test set for evaluation, with one demonstration per class sampled from the training set. Experiments with more demonstrations yield similar outcomes (refer to Appendix F.1 for details). Results reflect averages from five random seeds.
@@
Figure 3: Relative sizes of $S_{wp}$, $S_{pq}$, and $S_{ww}$ in different layers on SST-2 and AGNews. Results of other datasets can be found in Appendix B. Initially, $S_{wp}$ occupies a significant proportion, but it gradually decays over layers, while $S_{pq}$ becomes the dominant one.
Figure 3 reveals that: (1) in shallow layers, $S_{pq}$, the significance of the information flow from label words to targeted positions, is low, while $S_{wp}$, the information flow from the text part to label words is high; (2) in deep layers, $S_{pq}$, the importance of information flow from label words to the targeted position becomes the dominant one. Notably, $S_{pq}$ and $S_{wp}$ usually surpass $S_{ww}$, suggesting that interactions involving label words outweigh others.
@@
Based on this, we propose the hypothesis that label words function as anchors in the ICL information flow. In shallow layers, label words gather information from demonstration examples to form semantic representations for deeper layers, while in deep layers, the model extracts the information from label words to form the final prediction. Figure 2 gives an illustration for our hypothesis.
@@ 2.2 Shallow Layers: Information Aggregation
In this part, we validate our hypothesis’ first component. We assume that the information aggregation in ICL relies on the information flow from the text part to label tokens, which is facilitated by the transformer’s attention mechanism. By manipulating the attention layer in the model to block this flow and examining the model behavior change, we validate the existence of the information aggregation process and its contribution to the final prediction.
@@
We retain the same test sample size of 1000 inputs as § 2.1. We use the same demonstration for a single random seed. To further validate our findings on larger models, we incorporate GPT-J (6B) (Wang and Komatsuzaki, 2021) in experiments, which exceeds GPT2-XL in model size and capacity.
@@
To block the information flow to label words, we isolate label words by manipulating the attention matrix $A$. Specifically, we set $A_{l}(p,i)(i<p)$ to 0 in the attention matrix $A_{l}$ of the $l$-th layer, where $p$ represents label words and $i$ represents preceding words. Consequently, in the $l$-th layer, label words cannot access information from the prior demonstration text.
@@
We use the following metrics to assess the impact of blocking information flow from the text part to label tokens: (1) Label Loyalty: measures the consistency of output labels with and without isolation. (2) Word Loyalty: employs the Jaccard similarity to compare the top-5 predicted words with and without isolation, capturing more subtle model output alterations (See Appendix C for details). Low loyalty indicates a profound impact of isolation on model predictions.
Figure 4: The impact of isolating label words versus randomly isolating non-label words within the first or last 5 layers. Isolating label words within the first 5 layers exerts the most substantial impact, highlighting the importance of shallow-layer information aggregation via label words.
@@
Figure 4 illustrates a notable influence on the model’s behavior when label words are isolated within the first 5 layers. Yet, this influence becomes inconsequential within the last 5 layers, or when random non-label words are used. This observation underlines the fundamental importance of shallow-layer information aggregation via label words in ICL. It also emphasizes the superiority of label words over non-label words. Further tests with variable numbers of layers reaffirm these findings (Appendix D). Moreover, similar results were obtained when testing ICL with semantically unrelated labels (refer to Appendix F.2).
@@ 2.3 Deep Layers: Information Extraction
We proceed to validate the latter part of our hypothesis that the model extracts information from label words to form the final prediction. We denote the sum of the attention matrices in the $l$-th layer as $A_{l}$. [^4] Here we sum up the attention matrices of all attention heads in the $l$th layer for convenience of analysis. In deeper layers, we find a strong correlation between the attention distributions on the label words of the target position, represented as $(A_{l}(q,p_{1}),...,A_{l}(q,p_{C}))$, and the model’s final prediction, affirming our hypothesis. The experimental setup mirrors that discussed in § 2.2.
@@ 2.3.1 Experiments
We utilize the AUC-ROC score to quantify the correlation between $A_{l}(q,p_{i})$ and model prediction, which we denote as $\text{AUCROC}_{l}$ for the $l$-th layer. We prefer the AUC-ROC metric due to two primary reasons: (1) $A_{l}(q,p_{i})$ might differ from the probability of the model outputting label $i$ by a constant factor. As Kobayashi et al. (2020) points out, attention should be multiplied by the norm of the key vector to yield ’more interpretable attention’. The AUC-ROC metric can implicitly account for these factors, thus allowing us to uncover the correlation more effectively. (2) The proportion of different labels output by the model may be unbalanced. Using the AUC-ROC metric can help mitigate this issue, reducing disturbances caused by class imbalance.
Considering the residual mechanism of transformers, we can view each layer’s hidden state as the cumulative effect of all prior layer calculations. To quantify the accumulated contribution of the first $l$ layers to model prediction, we introduce $R_{l}$:
@@
This measure tracks the positive contribution above a baseline AUC-ROC threshold of 0.5. The value of $R_{l}$ signifies the proportional contribution of the first $l$ layers to the model prediction.